Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2001-293_2001-299

Mission Operations Days: 2001/293 to 2001/299
October 20, 2001 00:00:00 GMT to October 26, 2001 00:00:00 GMT


Data Loss on DOY 292

See “Non-MODIS Significant Events” below for details.


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - A Side
ComponentState(s)Comment
BlackbodyA On (290 K); B OffNominal
Calibration ElectronicsA On; B OffNominal
Control ProcessorA On; B OffNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA On; B OffNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 On; #3 & #4 OffNominal
Format ProcessorA On; B OffNominal
Power Supplies#1 On; #2 OffNominal
PV VIS FPAA On; B OffNominal
PV NIR FPAA On; B OffNominal
PV SMIR FPAA On; B OffNominal
PV LWIR FPAA On; B OffNominal
PC LWIR FPAA On; B OffNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA On; B OffNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA Off; B OnNominal
Flight SoftwareRev BD + 3 set of patches
  • FR Error Counting and Statistics (2 patches)
  • FR Band 5 Gain Change (20 commands)
  • FR BB Ideal Output Values (1 patch)
Nominal
Inhibit Ids SetNoneNominal
TMON 66 (SRCA A)DisabledNominal
TMON 67 (SRCA B)DisabledNominal
TMON 68 (FIFO)EnabledNominal
TMON 69 (EPS Current)DisabledNominal

This Week's Completed MODIS Activities:

Saturday, October 20, 2001

2001/293 05:00:00 ATC OA26: BB Temperature Stepping (BB to 290K)

Sunday, October 21, 2001

None

Monday, October 22, 2001

None

Tuesday, October 23, 2001

None

Wednesday, October 24, 2001

2001/297 21:14:48-21:26:16 ATC OA15: SD/SDSM Open
2001/297 22:53:41-23:05:09 ATC OA16: SD/SDSM Screened

Thursday, October 25, 2001

None

Friday, October 26, 2001

None

This Week's Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

Two SD/SDSM calibrations will be scheduled during the upcoming week.

Maneuvers:

The next Drag Makeup Maneuver will occur on 2001/303 (October 30, 2001) at 19:32:00.

The next Lunar Calibration roll maneuver is scheduled for day 2001/309 (November 5, 2001) with data collection from 18:15:11 to 18:21:41 (Space Craft roll time of 18:09:07 to 18:26:50). The roll angle will be -11.8237 degrees.

The topics of Terra's orbit inclination adjust maneuvers and eventual equator crossing time has resurfaced. Terra's equator crossing mean local time (MLT) requirement is 10:30 +/- 00:15. The current crossing time is about 10:35 and decreasing by 1-2 seconds per day. The orbit management talks have begun because, under current conditions, Terra will approach the SAC-C spacecraft very closely in April of 2002. Further details will follow as a finalized plan emerges.

MODIS Anomalies:

Formatter Error Events:

As of 2001/299, 15:34:30, the FR error counter (MOD_SS_CP_STATUS_08) has “rolled over” 43 times since the A-Side restart following the PS-2 Anomaly, and 41 times since the latest string of errors started on August 15. This is a 16-bit signed counter, therefore it “rolls over” every 32,768 counts. At this time, the total number of errors seen is roughly 1.4 million.

The FR Error Rate roll over period appears to be asymptotically approaching roughly an 11-hour period. The period decreases and increases based on cooling or heating (respectively) of the MODIS electronics from internal calibrations (BB, SRCA, and SD/SDSM) activities. However, once an activity has completed, and the MODIS electronic thermal balance returns to “non- activity” levels, the error rate returns to it previous 11-hour roll over period.

An error tracking spreadsheets have been updated and can be retrieved from:

ftp://www.mcst.ssai.biz/pub/permanent/IOT/am1/analysis/FR_events/

An MS Excel spreadsheet called, AM1_FR_error_rollovers_MMDDYY.xls contains the data for the times each of the error count “roll over” occurs (with the MMDDYY denoting the date the spreadsheet was made on, the more recent date containing the most up-to-date information).

General Instrument Comments:

MODIS is in Science Mode on the A-side with the SVD and NAD open, operating with the Formatter Resolution Flight Software Patches (FRFSP).

Note: SD/SDSM calibrations were planned to be reduced to once per two weeks. This plan has been revised; SD/SDSM calibrations will continue to be scheduled once each week until the end of November.

MODIS Telemetry Trends:

Instrument temperatures across the board are rising slightly as of late, although the orbit beta angle has not yet reached its minimum.

Non-MODIS Significant Events:

The Network Communications Center (NCC) has been rejecting and/or canceling Terra TDRSS contact times (likely to support DOD satellites). This has resulted in at least one incidence of data loss.

On DOY 2001/292, at 18:11:37 there was a data loss due to contact scheduling problems (the reason described above). This resulted in 1,702,006 CADUs of MODIS Data being lost. (A CADU is 7168 bits.) MODIS was in FR Day Rate at this time, and the data corresponds to roughly 19:17 (MM:SS) of Day Rate Science Data.

The full Data Loss Report can be viewed at:

http://jupiter02.gsfc.nasa.gov/

Select "Search Data Losses", and provide the following information:

Data Type: MOD Loss Start: 292/18:11:37

Then press the big SUBMIT button.

Limited Life Item Status:

SRCA 10W Lamp #1: 192.2 of 500 hours
SRCA 10W Lamp #2: 146.6 of 500 hours
SRCA 10W Lamp #3: 155.8 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 558.3 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1739 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements